Hi A car travels 200km the 2 highest speed of a car is 80 km/h the journey take two hours what is the average speed of the car?



1
Expert's answer
2021-01-12T12:10:24-0500

"s=200(km)"


"t=2(hours)"


"v_{avr}=\frac{s}{t}=\frac{200}{2}=100(km/h)" . Answer